# Joint Cooperative and Non-Cooperative Localization in WSNs with Distributed Scaled Proximal ADMM Algorithms

### v1

@@ 2. Using the update rule for $\bm{\lambda}_{i}^{t+1}$, we obtain:
The first inequality follows from Lemma A.8.1, and the second from equation (C.1). Hence, for all $i\in\mathcal{N}$, the sequence $\{\|\mathbf{A}_{i}\mathbf{z}_{i}^{t+1}\|\}_{t\geq 1}$ is bounded. Similarly, applying Lemma A.8.1 and equation (C.1) again, we find that $\{\|\mathbf{z}_{i}^{t+1}-\mathbf{z}_{i}^{t}\|\}_{t\geq 1}$ is also bounded. Since for all $i\in\mathcal{N}$:
$\mathbf{H}_{i}:$ $=\begin{bmatrix}\mathbf{1}_{N_{i}}&\mathbf{O}_{N_{i}}&-\mathbf{I}_{N_{i}}&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}\\ 1&\mathbf{0}_{N_{i}}^{T}&\mathbf{0}_{N_{i}}^{T}&-1&\mathbf{0}_{N_{i}}^{T}&\mathbf{0}_{N_{i}}^{T}\end{bmatrix}\otimes\mathbf{I}_{n}$
$\mathbf{A}_{i}:$ $=\begin{bmatrix}\mathbf{1}_{N_{i}}&-\mathbf{I}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}\\ \mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{1}_{N_{i}}&-\mathbf{I}_{N_{i}}&\mathbf{O}_{N_{i}}\\ \mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{1}_{N_{i}}&\mathbf{O}_{N_{i}}&-\mathbf{I}_{N_{i}}\end{bmatrix}\otimes\mathbf{I}_{n}$ (G.3)
we can derive through straightforward matrix operations that:
$\begin{bmatrix}\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&-\mathbf{I}_{N_{i}}&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}\\ 0&\mathbf{0}_{N_{i}}^{T}&\mathbf{0}_{N_{i}}^{T}&-1&\mathbf{1}_{N_{i}}^{T}&\mathbf{1}_{N_{i}}^{T}\end{bmatrix}\otimes\mathbf{I}_{n}\cdot\mathbf{H}_{i}^{T}$ $=\mathbf{I}_{(N_{i}+1)n},\quad$
$\begin{bmatrix}\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&-\mathbf{I}_{N_{i}}&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}\\ 0&\mathbf{0}_{N_{i}}^{T}&\mathbf{0}_{N_{i}}^{T}&-1&\mathbf{1}_{N_{i}}^{T}&\mathbf{1}_{N_{i}}^{T}\end{bmatrix}\otimes\mathbf{I}_{n}\cdot\mathbf{A}_{i}^{T}$ $=\mathbf{O}_{(N_{i}+1)n\times 3N_{i}n}.$ (G.4)
Moreover, substituting equation (III) into the update rule of $\tilde{\mathbf{z}}_{i}^{t+1}$ yields:
@@
Multiplying both sides of equation (G) by the matrix
$\begin{bmatrix}\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&-\mathbf{I}_{N_{i}}&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}\\ 0&\mathbf{0}_{N_{i}}^{T}&\mathbf{0}_{N_{i}}^{T}&-1&\mathbf{1}_{N_{i}}^{T}&\mathbf{1}_{N_{i}}^{T}\end{bmatrix}$
and then taking the norm, we obtain:
$\|\mathbf{H}_{i}\mathbf{z}_{i}^{t+1}\|$ $=\|\begin{bmatrix}\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&-\mathbf{I}_{N_{i}}&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}\\ 0&\mathbf{0}_{N_{i}}^{T}&\mathbf{0}_{N_{i}}^{T}&-1&\mathbf{1}_{N_{i}}^{T}&\mathbf{1}_{N_{i}}^{T}\end{bmatrix}\otimes\mathbf{I}_{n}\cdot\mathbf{U}_{i}(\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{z}_{i}^{t})+\mathbf{D}_{i}\mathbf{w}_{i}^{t}\|$
@@
where $d_{\max}=\max\left(\{d_{i,j}\mid i\in\mathcal{N},j\in\mathcal{N}_{i}\}\cup\{r_{i}\mid i\in\mathcal{N}\}\right)$. The first inequality follows from equation (G.4), and the second is based on the triangle inequality. Since the right-hand side of the inequality is bounded, it follows that $\{\|\mathbf{H}_{i}\mathbf{z}_{i}^{t}\|\}_{t\geq 1}$ is bounded for all $i\in\mathcal{N}$. Recalling the definitions $\mathbf{z}_{i}=[\mathbf{x}_{i}^{T},(\mathbf{p}_{i}^{-})^{T},(\mathbf{p}_{i}^{+})^{T},\mathbf{y}_{i}^{T},(\mathbf{q}_{i}^{-})^{T},(\mathbf{q}_{i}^{+})^{T}]^{T}$, along with $\mathbf{H}_{i}$ and $\mathbf{A}_{j}$, it follows that:
$\mathbf{H}_{i}\mathbf{z}_{i}^{t}$ $=\text{vec}\left(\mathbf{x}_{i}^{t}-(\mathbf{p}_{i,j}^{+})^{t}\mid j\in\mathcal{N}_{j},\;\mathbf{x}_{i}^{t}-\mathbf{y}_{i}^{t}\right),$
$\mathbf{A}_{j}\mathbf{z}_{j}^{t}$ $=\text{vec}\left(\mathbf{x}_{j}^{t}-(\mathbf{p}_{j,i}^{-})^{t}\mid i\in\mathcal{N}_{j},\mathbf{y}_{j}^{t}-(\mathbf{q}_{j,i}^{-})^{t}\mid i\in\mathcal{N}_{j},\mathbf{x}_{j}^{t}-(\mathbf{q}_{j,i}^{+})^{t}\mid i\in\mathcal{N}_{j}\right).$ (G.6)
@@
Since at least one anchor exists, combining equation (G) and equation (G) yields:
@@
Because both $\|\mathbf{H}_{i}^{t}\mathbf{z}_{i}\|$ and $\|\mathbf{A}_{j}^{t}\mathbf{z}_{j}\|$ are bounded for all $i\in\mathcal{N}$, it follows from equation (G.6) that the right-hand side of inequality (G) is bounded. Hence, the sequence $\{\|\mathbf{p}_{j}^{t}\|\}_{t\geq 1}$ is bounded for all $j\in\mathcal{N}_{i}$, $i\in\mathcal{A}$.
By a similar argument as in equations (G)–(G), we can also prove that for the neighbor sensor nodes $j\in\mathcal{N}_{i}\cap\mathcal{A}$, the sequence $\{\mathbf{x}_{i}^{t}\}_{t\geq 1}$ is bounded. Since the graph is connected, we conclude that the sequence $\{\mathbf{x}_{i}^{t}\}_{t\geq 1}$ is bounded for all $i\in\mathcal{N}$.
@@
$\|\mathbf{y}_{i}^{t}\|\leq\|\mathbf{x}_{i}^{t}\|+\|\mathbf{x}_{i}^{t}-\mathbf{y}_{i}^{t}\|,\forall i\in\mathcal{N},$
and the fact that $\|\mathbf{H}_{i}^{t}\mathbf{z}_{i}\|$ is bounded for all $i\in\mathcal{N}$, we conclude that the sequence $\{\mathbf{y}_{i}^{t}\}_{t\geq 1}$ is also bounded.
@@
$\|\mathbf{z}_{i}^{t}\|=$ $\|\mathbf{x}_{i}^{t}\|+\|\mathbf{y}_{i}^{t}\|+\sum_{j\in\mathcal{N}_{i}}(\|(\mathbf{p}_{i,j}^{-})^{t}\|+\|(\mathbf{p}_{i,j}^{+})^{t}\|+\|(\mathbf{q}_{i,j}^{-})^{t}\|+\|(\mathbf{q}_{i,j}^{+})^{t}\|)$
$\leq$ $2\|\mathbf{x}_{i}^{t}\|+\|\mathbf{x}_{i}^{t}-\mathbf{y}_{i}^{t}\|+\sum_{i\in\mathcal{N}_{j}}(2\|\mathbf{x}_{i}^{t}\|+\|\mathbf{x}_{i}^{t}-(\mathbf{p}_{i,j}^{-})^{t}\|+\|\mathbf{x}_{i}^{t}-(\mathbf{p}_{i,j}^{+})^{t}\|$
$+2\|\mathbf{y}_{i}^{t}\|+\|\mathbf{y}_{i}^{t}-(\mathbf{q}_{i,j}^{-})^{t}\|+\|\mathbf{y}_{i}^{t}-(\mathbf{q}_{i,j}^{+})^{t}\|),\quad\forall i\in\mathcal{N}.$ (G.10)
@@
3. Rearranging the terms in equation (G) and taking the norm yields:
$\|\bm{\lambda}_{i}^{t}\|\leq\|\mathbf{A}_{i}^{T}\bm{\lambda}_{i}^{t}\|\leq\|\mathbf{H}_{i}^{T}\mathbf{H}_{i}\mathbf{z}_{i}^{t}\|+\|W_{i}(\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{z}_{i}^{t})\|+\|\mathbf{H}_{i}^{T}\mathbf{D}_{i}\mathbf{w}_{i}^{t}\|+\|c\mathbf{A}_{i}^{T}\mathbf{A}_{i}\mathbf{z}_{i}^{t}\|.$
The first inequality above holds because equation (G) indicates that the smallest eigenvalue of $\mathbf{A}_{i}\mathbf{A}_{i}^{T}$ is 1. The second inequality follows from the triangle inequality. Since the right-hand side of the second inequality in equation (G) is bounded, it follows that the sequence $\{\bm{\lambda}_{i}^{t}\}_{t\geq 1}$ is bounded for all $i\in\mathcal{N}$.
@@ 4. By the triangle inequality, we have
$\|\bm{\xi}^{t+1}\|\leq$ $\sum_{i\in\mathcal{N}}\alpha_{i}\|\mathbf{z}_{i}^{t+1}-\mathbf{z}_{i}^{t}\|+\sum_{i\in\mathcal{N}}\left(\rho+d_{\max}\sqrt{N_{i}+2}\right)\|\mathbf{w}_{i}^{t+1}-\mathbf{w}_{i}^{t}\|$
@@
$\alpha_{i}:=$ $c(2N_{i}+1)+N_{i}+2+2(\kappa_{1}+\kappa_{2})(1+2c)(2+N_{\max})$
$+\frac{2[(2c+1)N_{i}+1](c+1)\sqrt{3nN_{\text{sum}}(2N_{i}+1)}}{\sqrt{\tilde{\tau}_{\min}}},$
@@
$C_{2}:=$ $\sqrt{3N}\cdot\max\{\alpha_{i},\rho+d_{\max}\sqrt{N_{i}+2},\beta_{i},i\in\mathcal{N}\}.$ (G.14)
@@
where the definitions of $\bm{\xi}_{\tilde{\mathbf{z}}_{i}^{t+1}},\bm{\xi}_{\mathbf{w}_{i}^{t}},\bm{\xi}_{\mathbf{z}_{i}^{t}}$ can be found in equation (G). Applying the update rule of $\tilde{\mathbf{z}}_{i}^{t+1}$ yields
@@
Combining equations (G.16) and (G) gives $\bm{\xi}_{\mathbf{z}_{i}^{t+1}}=\nabla_{\mathbf{z}_{i}^{t+1}}\mathcal{P}^{t+1}$. According to the optimality condition of problem (26b) with respect to variable $\mathbf{v}$, we have
@@
Combining equations (G.17) and (G) shows that $\bm{\xi}_{\mathbf{z}_{i}^{t+1}}\in\partial_{\mathbf{z}_{i}^{t+1}}\mathcal{P}^{t+1}$. Using the update rule of $\bm{\lambda}_{i}^{t+1}$, we have
@@
where $\sigma_{\max}(\mathbf{H}_{i}\mathbf{H}_{i}^{T})$ and $\sigma_{\max}(\mathbf{A}_{i}\mathbf{A}_{i}^{T})$ denote the largest eigenvalues of $\mathbf{H}_{i}\mathbf{H}_{i}^{T}$ and $\mathbf{A}_{i}\mathbf{A}_{i}^{T}$, respectively. Taking square roots of inequality (B) and using the fact that $\|\mathbf{x}\|_{2}\leq\|\mathbf{x}\|_{1}\leq\sqrt{N}\|\mathbf{x}\|_{2}$ for any $\mathbf{x}\in\mathbb{R}^{N}$, we get
@@
Moreover, from equations (III) and (G) we obtain
$\|\mathbf{U}_{i}\|=\max\big\{2[(c+1)N_{i}+1],2(2cN_{i}+1)\big\}\leq 2[(2c+1)N_{i}+1],$
$\sigma_{\max}(\mathbf{A}_{i}\mathbf{A}_{i}^{T})=2N_{i}+1,\quad\sigma_{\max}(\mathbf{H}_{i}\mathbf{H}_{i}^{T})=N_{i}+2.$ (G.20)
Substituting equations (G), (G), and (G.20) into equation (G.18), we obtain
$\|\bm{\xi}^{t+1}\|\leq$ $\sum_{i\in\mathcal{N}}\alpha_{i}\|\mathbf{z}_{i}^{t+1}-\mathbf{z}_{i}^{t}\|+\sum_{i\in\mathcal{N}}\left(\rho+d_{\max}\sqrt{N_{i}+2}\right)\|\mathbf{w}_{i}^{t+1}-\mathbf{w}_{i}^{t}\|$

### v2

@@ 2. Using the update rule for $\bm{\lambda}_{i}^{t+1}$, we obtain:
The first inequality follows from Lemma A.8.1, and the second from equation (C.1). Hence, for all $i\in\mathcal{N}$, the sequence $\{\|\mathbf{A}_{i}\mathbf{z}_{i}^{t+1}\|\}_{t\geq 1}$ is bounded. Similarly, applying Lemma A.8.1 and equation (C.1) again, we find that $\{\|\mathbf{z}_{i}^{t+1}-\mathbf{z}_{i}^{t}\|\}_{t\geq 1}$ is also bounded.
We construct a matrix $\mathbf{B}_{i}$ of the form
$\mathbf{B}_{i}$ $:=\begin{bmatrix}\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&-\mathbf{I}_{N_{i}}&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}&...&\mathbf{0}_{N_{i}}&\mathbf{O}_{N_{i}}&\mathbf{O}_{N_{i}}\\ 0&\mathbf{0}_{N_{i}}^{T}&\mathbf{0}_{N_{i}}^{T}&-1&\mathbf{1}_{N_{i}}^{T}&\mathbf{1}_{N_{i}}^{T}&...&-1&\mathbf{1}_{N_{i}}^{T}&\mathbf{1}_{N_{i}}^{T}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&&\vdots&\vdots&\vdots\\ 0&\mathbf{0}_{N_{i}}^{T}&\mathbf{0}_{N_{i}}^{T}&-1&\mathbf{1}_{N_{i}}^{T}&\mathbf{1}_{N_{i}}^{T}&...&-1&\mathbf{1}_{N_{i}}^{T}&\mathbf{1}_{N_{i}}^{T}\end{bmatrix}\otimes\mathbf{I}_{n}\in\mathbb{R}^{(N_{i}+M)n\times(1+M)(1+2N_{i})n}$ (G.3)
This matrix is designed to satisfy the condition, for all $i\in\mathcal{N}$:
$\mathbf{B}_{i}\cdot\mathbf{H}_{i}^{T}\mathbf{H}_{i}$ $=\mathbf{H}_{i},\quad$
$\mathbf{B}_{i}\cdot\mathbf{A}_{i}^{T}\mathbf{A}_{i}$ $=\mathbf{O}_{(N_{i}+M)n\times(1+2N_{i})n}.$ (G.4)
Moreover, substituting equation (3) into the update rule of $\tilde{\mathbf{z}}_{i}^{t+1}$ yields:
@@
[…]
$\|\mathbf{H}_{i}\mathbf{z}_{i}^{t+1}\|$ $=\|\mathbf{B}_{i}\cdot\mathbf{U}_{i}(\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{z}_{i}^{t})+\mathbf{D}_{i}\mathbf{w}_{i}^{t}\|$
@@
where $\color[rgb]{0,0,0}d_{\max}=\max\left(\{d_{i,j}\mid i\in\mathcal{N},j\in\mathcal{N}_{i}\}\cup\{r_{i,{\color[rgb]{0,0,0}k}}\mid i\in\mathcal{N},k\in\mathcal{M}_{i}\}\right)$}. The first inequality follows from equation (G.4), and the second is based on the triangle inequality. Since the right-hand side of the inequality is bounded, it follows that $\{\|\mathbf{H}_{i}\mathbf{z}_{i}^{t}\|\}_{t\geq 1}$ is bounded for all $i\in\mathcal{N}$. Recalling the definitions $\mathbf{z}_{i}$, along with $\mathbf{H}_{i}$ and $\mathbf{A}_{j}$, it follows that:
$\mathbf{H}_{i}\mathbf{z}_{i}^{t}$ $=\text{vec}\left(\mathbf{x}_{i}^{t}-(\mathbf{p}_{i,j}^{+})^{t}\mid j\in\mathcal{N}_{j}\quad,\quad\mathbf{x}_{i}^{t}-\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}\mid{\color[rgb]{0,0,0}k\in\mathcal{M}_{i}}\right),$
$\mathbf{A}_{j}\mathbf{z}_{j}^{t}$ $=\text{vec}\left(\mathbf{x}_{j}^{t}-(\mathbf{p}_{j,i}^{-})^{t}\mid i\in\mathcal{N}_{j}\quad,\quad\mathbf{y}_{j}^{t}-(\mathbf{q}_{j,i,{\color[rgb]{0,0,0}k}}^{-})^{t}\mid i\in\mathcal{N}_{j},k\in\mathcal{M}\quad,\quad\mathbf{x}_{j}^{t}-(\mathbf{q}_{j,i,{\color[rgb]{0,0,0}k}}^{+})^{t}\mid i\in\mathcal{N}_{j},k\in\mathcal{M}\right).$ (G.6)
@@
Since at least one anchor exists, combining equation (A.7) and equation (A.7) yields:
@@
Because both $\|\mathbf{H}_{i}^{t}\mathbf{z}_{i}\|$ and $\|\mathbf{A}_{j}^{t}\mathbf{z}_{j}\|$ are bounded for all $i\in\mathcal{N}$, it follows from equation (G.6) that the right-hand side of inequality (A.7) is bounded. Hence, the sequence $\{\|\mathbf{p}_{j}^{t}\|\}_{t\geq 1}$ is bounded for all $j\in\mathcal{N}_{i}$, $i\in\mathcal{A}$.
By a similar argument as in equations (A.7)–(A.7), we can also prove that for the neighbor sensor nodes $j\in\mathcal{N}_{i}\cap\mathcal{A}$, the sequence $\{\mathbf{x}_{i}^{t}\}_{t\geq 1}$ is bounded. Since the graph is connected, we conclude that the sequence $\{\mathbf{x}_{i}^{t}\}_{t\geq 1}$ is bounded for all $i\in\mathcal{N}$.
@@
$\|\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}\|\leq\|\mathbf{x}_{i}^{t}\|+\|\mathbf{x}_{i}^{t}-\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}\|,\forall k\in\mathcal{M},i\in\mathcal{N},$
and the fact that $\|\mathbf{H}_{i}^{t}\mathbf{z}_{i}\|$ is bounded for all $i\in\mathcal{N}$, we conclude that the sequence $\{\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}\}_{t\geq 1}$ is also bounded.
@@
$\|\mathbf{z}_{i}^{t}\|=$ $\|\mathbf{x}_{i}^{t}\|+\|\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}\|+\sum_{j\in\mathcal{N}_{i}}(\|(\mathbf{p}_{i,j}^{-})^{t}\|+\|(\mathbf{p}_{i,j}^{+})^{t}\|+\|(\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{-})^{t}\|+\|(\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{+})^{t}\|)$
$\leq$ $2\|\mathbf{x}_{i}^{t}\|+\sum_{k\in\mathcal{M}_{i}}\|\mathbf{x}_{i}^{t}-\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}\|+\sum_{i\in\mathcal{N}_{j}}\Big(2\|\mathbf{x}_{i}^{t}\|+\|\mathbf{x}_{i}^{t}-(\mathbf{p}_{i,j}^{-})^{t}\|+\|\mathbf{x}_{i}^{t}-(\mathbf{p}_{i,j}^{+})^{t}\|$
$+\sum_{k\in\mathcal{M}}(2\|\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}\|+\|\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}-(\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{-})^{t}\|+\|\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}-(\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{+})^{t}\|)\Big),\quad\forall i\in\mathcal{N}.$ (G.10)
@@
3. Rearranging the terms in equation (A.7) and taking the norm yields:
$\|\bm{\lambda}_{i}^{t}\|\leq\|\mathbf{A}_{i}^{T}\bm{\lambda}_{i}^{t}\|\leq\|\mathbf{H}_{i}^{T}\mathbf{H}_{i}\mathbf{z}_{i}^{t}\|+\|U_{i}(\tilde{\mathbf{z}}_{i}^{t+1}-\mathbf{z}_{i}^{t})\|+\|\mathbf{H}_{i}^{T}\mathbf{D}_{i}\mathbf{w}_{i}^{t}\|+\|c\mathbf{A}_{i}^{T}\mathbf{A}_{i}\mathbf{z}_{i}^{t}\|.$
The first inequality above holds because the smallest eigenvalue of $\mathbf{A}_{i}\mathbf{A}_{i}^{T}$ is 1. The second inequality follows from the triangle inequality. Since the right-hand side of the second inequality in equation (A.7) is bounded, it follows that the sequence $\{\bm{\lambda}_{i}^{t}\}_{t\geq 1}$ is bounded for all $i\in\mathcal{N}$.
@@ 4. By the triangle inequality, we have
$\|\bm{\xi}^{t+1}\|\leq$ $\sum_{i\in\mathcal{N}}\alpha_{i}\|\mathbf{z}_{i}^{t+1}-\mathbf{z}_{i}^{t}\|+\sum_{i\in\mathcal{N}}\left(\rho+d_{\max}\sqrt{N_{i}+M_{i}+1}\right)\|\mathbf{w}_{i}^{t+1}-\mathbf{w}_{i}^{t}\|$
@@
$\alpha_{i}:=$ $c(2N_{i}+1)+N_{i}+M_{i}+1+2(\kappa_{1}+\kappa_{2})(1+c)(2+N_{\max})$
$+\frac{2[(c+1)N_{i}+1](c+1)\sqrt{3nN_{\text{sum}}(2N_{i}+1)}}{\sqrt{\tilde{\tau}_{\min}}},$
@@
$C_{2}:=$ $\sqrt{3N}\cdot\max\{\alpha_{i},\rho+d_{\max}\sqrt{N_{i}+M_{i}+1},\beta_{i},i\in\mathcal{N}\}.$ (G.14)
@@
where the definitions of $\bm{\xi}_{\tilde{\mathbf{z}}_{i}^{t+1}},\bm{\xi}_{\mathbf{w}_{i}^{t}},\bm{\xi}_{\mathbf{z}_{i}^{t}}$ can be found in equation (A.7). Applying the update rule of $\tilde{\mathbf{z}}_{i}^{t+1}$ yields
@@
Combining equations (G.16) and (A.7) gives $\bm{\xi}_{\mathbf{z}_{i}^{t+1}}=\nabla_{\mathbf{z}_{i}^{t+1}}\mathcal{P}^{t+1}$. According to the optimality condition of problem (25b) with respect to variable $\mathbf{v}$, we have
@@
Combining equations (G.17) and (A.7) shows that $\bm{\xi}_{\mathbf{z}_{i}^{t+1}}\in\partial_{\mathbf{z}_{i}^{t+1}}\mathcal{P}^{t+1}$. Using the update rule of $\bm{\lambda}_{i}^{t+1}$, we have
@@
where $\sigma_{\max}(\mathbf{H}_{i}\mathbf{H}_{i}^{T})$ and $\sigma_{\max}(\mathbf{A}_{i}\mathbf{A}_{i}^{T})$ denote the largest eigenvalues of $\mathbf{H}_{i}\mathbf{H}_{i}^{T}$ and $\mathbf{A}_{i}\mathbf{A}_{i}^{T}$, respectively. Taking square roots of inequality (A.2) and using the fact that $\|\mathbf{x}\|_{2}\leq\|\mathbf{x}\|_{1}\leq\sqrt{N}\|\mathbf{x}\|_{2}$ for any $\mathbf{x}\in\mathbb{R}^{N}$, we get
@@
Moreover, from equations (3) we obtain
$\|\mathbf{U}_{i}\|=\max\big\{2[(c+1)N_{i}+M_{i}],2(2cN_{i}+1)\big\}\leq 2[(2c+1)N_{i}+M_{i}+1],$
$\sigma_{\max}(\mathbf{A}_{i}\mathbf{A}_{i}^{T})=2N_{i}+1,\quad\sigma_{\max}(\mathbf{H}_{i}\mathbf{H}_{i}^{T})=N_{i}+M_{i}+1.$ (G.20)
Substituting equations (A.7), (A.7), and (G.20) into equation (G.18), we obtain
$\|\bm{\xi}^{t+1}\|\leq$ $\sum_{i\in\mathcal{N}}\alpha_{i}\|\mathbf{z}_{i}^{t+1}-\mathbf{z}_{i}^{t}\|+\sum_{i\in\mathcal{N}}\left(\rho+d_{\max}\sqrt{N_{i}+M_{i}+1}\right)\|\mathbf{w}_{i}^{t+1}-\mathbf{w}_{i}^{t}\|$